### v2

# Zhiwei Zuo Zhuo Tang Kenli Li Anwitaman Datta

Abstract
Machine unlearning techniques, which involve retracting data records and reducing influence of said data on trained models, help with the user privacy protection objective but incur significant computational costs. Weight perturbation-based unlearning is a general approach, but it typically involves globally modifying the parameters. We propose fine-grained Top-K and Random-k parameters perturbed inexact machine unlearning strategies that address the privacy needs while keeping the computational costs tractable.
In order to demonstrate the efficacy of our strategies we also tackle the challenge of evaluating the effectiveness of machine unlearning by considering the model’s generalization performance across both unlearning and remaining data. To better assess the unlearning effect and model generalization, we propose novel metrics, namely, the forgetting rate and memory retention rate. However, for inexact machine unlearning, current metrics are inadequate in quantifying the degree of forgetting that occurs after unlearning strategies are applied. To address this, we introduce SPD-GAN, which subtly perturbs the distribution of data targeted for unlearning. Then, we evaluate the degree of unlearning by measuring the performance difference of the models on the perturbed unlearning data before and after the unlearning process. By implementing these innovative techniques and metrics, we achieve computationally efficacious privacy protection in machine learning applications without significant sacrifice of model performance. Furthermore, this approach provides a novel method for evaluating the degree of unlearning.
keywords:
Machine Unlearning, Perturbation, User Privacy, Forgetting
[…]
1 Introduction
Users may want the removal of personal data in possession of an organization. In several jurisdictions, this has compliance implications under regulations such as the General Data Protection Regulation (GDPR) in the European Union goddard2017eu and the California Consumer Privacy Act in the United States goldman2020introduction . Machine learning models trained using such data retain their influence, thus risking privacy compromises cheng2023privacy ; wu2023privacy . For instance, in a recommendation system using collaborative filtering schafer2007collaborative ; isinkaye2015recommendation , if the model does not eliminate the influence of removed data (which we call ‘unlearning data’), it may continue to make recommendations based on user similarity.
Separately, one may want to remove the influence of a subset of data from a model trained with a corpus of data, possibly because the information in the subset corpus is obsolete, or even wrong e.g., determined to be fake/misinformation subsequent to its original use in model training.
The task to address this concern by eliminating the influence of unlearning data on trained models has thus emerged in the recent years as the budding topic of machine unlearning.
Retraining from scratch, due to its extensive time and computational resource consumption, has become impractical, thus making machine unlearning extremely necessary. Common unlearning strategies can be divided into two catagories: exact unlearning and inexact unlearning.
Exact unlearning thudi2022necessity ; bourtoule2021machine aim to completely eliminate the influence of unlearning data on a model. Some researches try to evaluate the impact of data on models, the Influence function offers insights but faces limitations in deep learning contexts koh2017understanding ; basu2020influence ; li2023selective . In contrast, the Single-Step Sample Erasure (SSSE) peste2021ssse utilizes the Fisher Information Matrix to sidestep computational hurdles. In contrast, inexact machine unlearning DBLP:journals/corr/abs-2201-06640 ; suriyakumar2022algorithms ; neel2021descent aims to reduce time and resources requirements by selectively eliminating the data’s influence on the model. Current inexact machine unlearning methods based on model weight perturbation typically involve adding random Gaussian noise to all model parameters. We want to conduct unlearning using a more fine-gained perturbation. To that end, we design two strategies: Random-k and Top-K. By selectively perturbing a small subset of parameters, we aim to achieve the desired effect of machine unlearning efficiently while also reducing the impact on the overall model performance.
In the area of machine unlearning, model indistinguishability golatkar2020eternal ; kurmanji2023towards is crucial to consider. When different algorithms learn from the same data set, they may acquire very similar knowledge and features. Consequently, when faced with independent and identically distributed (i.i.d) datasets, even if an unlearning approach has eliminated the influence of unlearning data, the model may still show approximate accuracy on the unlearning data to the remaining data due to the model’s generalization property nakkiran2020distributional . This makes it challenging to evaluate the effectiveness of the unlearning process.
Consequently, to measure the effectiveness of unlearning, metrics can be divided into two aspects. The first is general performance evaluation, which includes assessing the accuracy marchant2022hard of unlearning model predictions, unlearn time corresponding to the unlearning request bourtoule2021machine ; becker2018organizational , relearn time for reaching the accuracy of source model tarun2021fast , membership inference attack (MIA) shokri2017membership ; hu2022membership . The second focuses on model indistinguishability, completeness, activation distance wu2020deltagrad , Jensen–Shannon (JS)-Divergence chundawat2023zero , and other metrics that quantify the differences between the unlearning and retraining models.
In the fine-grained inexact machine unlearning strategies we employ, we consider both the impact of model unlearning and the generalization performance of the model. A metric called forgetting rate is designed to assess the effect of model unlearning, while memory retention rate and similarity are employed to evaluate the model’s indistinguishability.
Figure 1: Random-k unlearning on CIFAR-10, accuracy of $D_{RE}$ and $D_{UL}$ after training for 50 epochs. Baseline is the test accuracy on the source model.
However, the existing metrics are unable to reveal the extent of unlearning. Suppose splitting training data into two parts: unlearning data $D_{UL}$ and remaining data $D_{RE}$ $\left(D=D_{UL}\cup D_{RE},len(D_{UL})\ll len(D_{RE})\right)$. As shown in Fig.1, while the accuracy may vary across $D_{UL}$ and $D_{RE}$ for each model, it consistently maintains a high accuracy for $D_{UL}$. Consequently, it becomes challenging to quantify the extent to which the model eliminates the influence of $D_{UL}$. Since the performance of unlearning model $M_{UL}$ on $D_{UL}$ and $D_{RE}$ is relatively small, in order to amplify the performance difference between the unlearning model $M_{UL}$ on $D_{UL}$ and $D_{RE}$, we propose a SPD-GAN (Slightly Perturbed Distribution-Generative Adversarial Network) to break the i.i.d attribute of the unlearning data. For the unlearning data $D_{UL}$, we introduce perturbations to it’s distribution, thus we gain perturbed unlearning data $D_{p}$. We limit the perturbations to be very slight so that the accuracy of source model $M$ on $D_{p}$ is close to the accuracy of $M$ on $D_{UL}$. Suppose $P(M,D)$ represents the performance of model $M$ on data $D$, seeing in the Equation.1, if the perturbed unlearning data $D_{p}$ exhibits significant performance difference between source model and unlearning model, we can then obtain the unlearning degree of each unlearning strategies.
$\centering P(M,D_{p})\gg P(M_{UL},D_{p})\quad s.t.\quad P(M,D_{p})\approx P(M,D_{UL})\@add@centering$ (1)
Our main contributions are:
1. We adopt two fine-grained inexact machine unlearning strategies. We propose strategies of Random-k and Top-K to quickly achieve the unlearning effect with minimal model parameters perturbed while ensuring its generalization properties (Section 3.2).
2. We design measurement metrics forgetting rate, memory retention rate and similarity to assess unlearning effectiveness and the unlearning model’s generalization properties. We carry out a theoretical analysis in Section 3.4 of the acceleration of inexact machine unlearning strategies, and represent it as acceleration ratio in our experiments (as discussed in Section 4.3.3).
3. We design a novel unlearning degree evaluation method by proposing a SPD-GAN to slightly perturb the unlearning data, using the performance difference of the perturbed unlearning data on the source model and unlearning model to approximate the degree of unlearning (Section 3.3). Experimental results show that our proposed Top-K can achieve the deepest degree of unlearning in comparison to other inexact unlearning strategies (as discussed in Section 4.3.5).
2 Related Works
2.1 User Privacy
Several important existing methods aim to protect user privacy. Traditional approaches focus on decoupling and generalization to eliminate user identifiers or use anonymization techniques bayardo2005data to generalize or compress user information. However, these methods may not be effective when prior knowledge is available.
Differential privacy (DP) dwork2006calibrating is a strict and provable privacy protection technique that adds noise in the stochastic gradient descent(SGD) process. Other approaches, such as dPA kocher2011introduction , add perturbation to the objective function, while PATE papernot2018scalable uses knowledge aggregation and transmission. However, the process of adding noise in DP is irreversible, making it unsuitable for machine unlearning tasks since the protection cannot be removed once data is protected by noise.
Federated learning (FL) is another method to protect user privacy, where training data is distributed among multiple devices and servers, and the model is trained in a decentralized manner without centralizing the data wei2020federated . FL methods based on differential privacy have been proposed to protect user privacy, with the principal focus on attacks xiao2022sca and defenses xiao2022sbpa in FL systems.
2.2 Machine Unlearning
Several approaches have emerged to address the limitations of retraining models from scratch for machine unlearning, since it is generally impractical. The typical approach is SISA (Sharded, Isolated, Sliced, and Aggregated) training, which splits data into shards and trains them in isolation for speed and effectiveness when unlearning. liu2022continual combines continual learning with private learning based on the SISA idea, but SISA-based methods often consume a lot of storage.
Other approaches focus on evaluating the impact of specific data points or sub-datasets on trained models. Influence function koh2017understanding approximates the parameter change when data is up-weighted by a small value, but its effectiveness in deep learning is limited by non-convex loss functions and Taylor’s approximation errors basu2020influence . Peste et al. propose a Single-Step Sample Erasure (SSSE) using the Fisher Information Matrix (FIM) to avoid expensive computation and Hessian inversion peste2021ssse . Some works use influence functions in Graph Neural Networks (GNNs) for unlearning, but these models are relatively shallow, and Hessian inversion is more accurate, allowing for exact unlearning.
Many model-related methods have been developed for machine unlearning. Ganhor et al. ganhor2022unlearning incorporate adversarial training into multVAE liang2018variational to remove implicit information of protected attributes. Liu et al. liu2022right study unlearning in federated learning and propose a fast retraining method to erase the impact of data samples from trained models. Some works discuss unlearning strategies in Bayesian methods rawat2022challenges , while others focus on random forests liu2021revfrf .
2.3 Adversarial Learning
Adversarial learning is an adversarial-based learning method that aims to enhance the robustness of machine learning algorithms against attacks. However, it may also lead to user privacy leakage, which can have severe consequences in areas such as biometric authentication.
The core concept of adversarial learning involves introducing minor perturbations to input data in order to cause incorrect classification. Szegedy et al. szegedy2013intriguing reveal the vulnerability of neural networks to adversarial examples and introduce the L-BFGS attack method. Due to the high time consumption associated with L-BFGS, Goodfellow et al. goodfellow2014explaining propose the fast gradient sign method (FGSM) which optimizes $L_{\infty}$distance. This method is later refined by Carlini and Wagner into the $C\&W$ method carlini2017towards . Additionally, Papernot et al. papernot2016limitations propose the Jacobian-based saliency map attack (JSMA), which optimizes under $L_{0}$ distance.
The concept of perturbation through the addition of noise is also employed in machine unlearning tasks. neel2021descent utilizes random Gaussian noise perturbations to the source model $M$’s parameters, constrained by ($\epsilon,\delta$), ensuring that the model remains indistinguishable from any point within a small neighborhood of the optimal model.
$\mathbb{P}(f_{m}(M(D))\in S)\leq e^{\epsilon}\mathbb{P}(f_{m}(M_{UL}(D_{RE}))\in S)+\delta$ (2)
$and\quad\mathbb{P}(f_{m}(M_{UL}(D_{RE}))\in S)\leq e^{\epsilon}\mathbb{P}(f_{m}(M(D))\in S)+\delta$
Where $f_{m}$ denotes a random mechanism, with model parameters as its input. Additionally, $S$ represents the subset of the model parameter space. In the status quo, existing adversary-based unlearning methods concentrate on ($\epsilon,\delta$)-perturbations, which are based on perturbing all parameters of the model.
3 Methodology
3.1 Problem Definition
In this paper, we focus on machine unlearning problem. Let $D$ represent the training data set which contains $n$ samples $\left(x_{1},y_{1}\right),\left(x_{2},y_{2}\right),...,\left(x_{n},y_{n}\right)$. For each sample, $x\in X$ and $y\in Y$. These data are posted by multiple users and collected by some institutions or organization and then labeled. Suppose $D=D_{UL}\cup D_{RE}$ and $D_{UL}\cap D_{RE}=\emptyset$, here $D_{RE}$ represents the remaining data after unlearning data removal. Given a trained model $M$ with parameter $w$ on data set $D$, when conducting machine unlearning, our goal is to gain a target model $M_{UL}$ with parameter $\theta$, which eliminates the influence of $D_{UL}$ and holds a quite good performance on $D_{RE}$.
We assume the posterior distribution of $M$ and $M_{UL}$ are $M\left(D_{UL}\right)\sim P_{1}$ and $M_{UL}\left(D_{UL}\right)\sim P_{2}$. When $P_{1}$ and $P_{2}$ are quite similar, we claim machine unlearning is conducted.
3.2 Our Approach
The parameters in the machine learning model store the high-dimensional features of the data, so in order to eliminate the influence of revoked data on the model, inspired by adversary learning szegedy2013intriguing , we consider adding slight perturbation to trained model and then perform several training epochs. Current machine unlearning methods based on parameters perturbation basically change all the parameters of the model or use tricks such as regularization or dropout to improve the efficiency of unlearning. We desire to perturb the parameters of the model at a finer granularity.
Not all parameters affect the model equally. The impact of the former layers on the model is mainly to extract the features of the data, while the later layers use these features to perform tasks. To this extent, it can be considered that perturbing the last few layers can achieve the purpose of inexact unlearning. However, only modifying the last few layers also means a lower degree of machine unlearning, as shown in Tab.2, where the accuracy $Acc_{UL}$ on $D_{UL}$ for EU-5, EU-10, CF-5, CF-10 is significantly higher than other unlearning strategies. In order to make the perturbation ratio decrease and make the perturbation more precise, we consider identifying several parameters in the model that have the greatest impact on performance to perturb, or to address the issue of reduced unlearning effect caused by perturbing the parameters of the last few layers. Thus, we propose two strategies for perturbing model parameters, namely Random-k and Top-K.
3.2.1 Parameter Selection and Perturbation
For Random-k, by setting the perturbation ratio $k$, randomly select the model parameters to be perturbed. For Top-K, we intend to calculate the parameter sensitivity on the data set to screen out the Top-K parameters that have the greatest impact on the model, so as to achieve the purpose of unlearning by performing parameter perturbation in a lighter way.
For the Top-K strategy, as previously mentioned, it is necessary to evaluate the sensitivity of each model parameter. Certainly, perturbing model parameters one by one can determine their sensitivity, but it’s time-consuming. To quickly obtain the sensitivity of parameters, we approximate it using the gradient. Specifically, after the model has converged, yielding the function $F(x,w)$, we introduce a small perturbation to the model parameters denoted as $\epsilon=\{\epsilon_{i,j}\}$. The change in the model’s output can be approximately measured using the following formula:
$\centering\bigtriangleup(x,w)=\sum_{i,j}\frac{\partial F(x,w)}{\partial w_{i,j}}*\epsilon_{i,j}\@add@centering$ (3)
Where $\bigtriangleup(x,w)$ represents the change in the model’s output due to the introduced perturbation $\epsilon$, and $\frac{\partial F(x,w)}{\partial\theta_{i,j}}=g_{i,j}(x)$ is the gradient of learned model to parameter $w$. This evaluation of sensitivity allows us to assess the impact of each parameter on the model’s predictions.
Focusing on aforementioned Equation 3, assuming the perturbation $\epsilon$ added to the model is the same constant, the calculation of sensitivity can be simplified as follows:
$\centering sensitivity_{i,j}=\|g_{i,j}(x)\|\@add@centering$ (4)
Figure 2: Machine Unlearning process using Random-k or Top-K perturbation strategies. When Random-k/Top-K strategies are applied, partial parameters changed (marked pink on unlearning model).
By adding such a constant perturbation across all model parameters, we can directly evaluate the sensitivity of each parameter’s contribution to the model’s overall output change. Only with one forward pass and back-propagation process, we obtain the sensitivity for each parameter.
Fig.2 illustrates the process of our approach. After completing the parameter screening, we introduce perturbations by drawing upon the insights from szegedy2013intriguing . Specifically, we apply a minute amount of noise to the selected parameters as follows:
$w_{i}=\epsilon*w_{rand},\quad w_{i}\in\{Top-K(w)\lor{Random-k(w)}\}$ (5)
Here $\epsilon$ is a coefficient with a small value. By introducing these perturbations, the model’s learned information is rendered inexact and less influential, allowing it to gradually fade out without abrupt removal. This approach offers a practical and robust solution for machine unlearning, enabling models to efficiently adapt to changing data while preserving essential knowledge from previous training.
3.2.2 Unlearning Process
As illustrated in Fig.2, small perturbations are introduced to the model $M$, followed by training the unlearning model $M_{UL}$ using $D_{RE}$ over multiple epochs to enable unlearning while still maintaining the performance of the remaining data. In this process, we gradually and iteratively adjust $M_{UL}$, nudging it towards a state that minimizes the influence of the unlearning data. This adaptive training strategy ensures that $M_{UL}$ retains its high predictive accuracy for the remaining data, while effectively ‘forgetting’ the data that is removed.
For the purpose of eliminating the influence of the data to be unlearned on the model, retraining from scratch, despite consuming substantial computational resources and time, can achieve the optimal unlearning effect. As far as the strategy for machine unlearning is concerned, we desire that the unlearning model $M_{UL}$ approximates the retraining model $M_{RE}$, thus ensuring that the unlearning is moving in the right direction. We aim for both the retraining model and the unlearning model to display similar performance when subjected to $D_{UL}$. Therefore, in the third step of unlearning, we input the data intended for unlearning into both models to obtain the corresponding output distributions.
Jensen-Shannon (JS) divergence is used to guide the model training of the unlearning process chundawat2023can . As a symmetric and always finite measure, JS divergence provides a reliable measure of the similarity or distance between two probability distributions. Suppose $x\in D_{UL}$, $P(w,x)$ and $P(\theta,x)$ are posterior distributions of retraining model $M_{RE}$ and unlearning model $M_{UL}$, respectively. $JS$ divergence is calculated as following:
$JS(P(\theta,x)||P(w,x))=\frac{1}{2}KL\left(P(\theta,x)||\frac{P(\theta,x)+P(w,x)}{2}\right)$ (6)
$+\frac{1}{2}KL\left(P(w,x)||\frac{P(\theta,x)+P(w,x)}{2}\right)$
where $KL(\cdot)$ is the Kullback-Leibler divergence. Since $P(\theta,x)$ and $P(w,x)$ are the output probability distributions of $M$ and $M_{UL}$ for $D_{UL}$, we have $\sum P(\theta,x)=\sum P(w,x)=1$, thus:
$JS(P(\theta,x)||P(w,x))=\frac{1}{2}\sum P(\theta,x)\log P(\theta,x)+$ (7)
$\begin{split}\frac{1}{2}\sum P(w,x)\log P(w,x)-\log(P(\theta,x)+P(w,x))-\log 2\end{split}$
In the context of machine unlearning, the JS divergence between distributions $P(w,x)$ and $P(\theta,x)$ can quantify the influence of $D_{UL}$ on $M_{UL}$. If the JS divergence is small, the influence of $D_{UL}$ on $M_{UL}$ is deemed to be minor. Therefore, during training, our goal should be to minimize the JS divergence between these distributions to ensure the effectiveness of unlearning. Moreover, given that the value of the JS divergence always falls between 0 and 1, we can intuitively understand the degree of similarity between the two distributions.
Mathematically, we articulate this as an optimization challenge. A novel loss function $L$ can be postulated, encompassing the original cross entropy loss in addition to a term associated with the JS divergence:
$L=\frac{1}{|D_{RE}|}\cdot\sum{y_{i,j}}\cdot log(p_{i,j})+\lambda\cdot JS(P(\theta,x)||P(w,x))$ (8)
Where $|D_{RE}|$ is the number of $D_{RE}$, $y_{i,j}$ is the true label of the $j-th$ category of the $i-th$ sample, and $p_{i,j}$ is the probability that the model predicts the $i-th$ sample as the $j-th$ category, $\lambda$ is used to balance the weight between the original loss and the JS divergence.
Algorithm 1 Top-K/Random-k
0:  Dataset $D$, unlearning data $D_{UL}$, remaining data $D_{RE}$, perturbed number $K$ / ratio $k$, coefficient $\epsilon$
1:  Train model $M$ on $D$
2:  $M_{UL}\leftarrow M$
3:  if Top-K then
4:     Pick a data sample $x$ in $D$, the gradient of $w$: $\frac{\partial F(x,w)}{\partial\theta_{i,j}}$
5:     for $w_{i}$ in $w$ do
6:        calculate sensitivity of each parameter:$sensitivity_{w_{i}}\leftarrow\|g_{w_{i}}(x)\|$
7:     end for
8:     $w_{j}\leftarrow\epsilon*w_{rand},w_{j}\in\{Top-K(w)\}$
9:  else
10:     $mask\leftarrow Rand(w.shape)$
11:     $mask\leftarrow(mask<k)*\epsilon$
12:     $w\leftarrow(1-mask)*w+mask*w_{rand}$
13:  end if
14:  while Epochs do
15:     $L_{1}\leftarrow\frac{1}{|D_{RE}|}\cdot\sum{y_{i,j}}\cdot log(p_{i,j})$
16:     Computes the JS divergence of the output distributions of $M$ and $M_{UL}$ for $D_{UL}$:
17:     $L_{2}\leftarrow JS(P(\theta,x)||P(w,x))$
18:     Minimize $L\leftarrow L_{1}+\lambda L_{2}$
19:  end while
Top-K and Random-k machine unlearning strategies are summarized in Algorithms 1.
3.3 Unlearning Effectiveness
By employing a fine-grained model parameter perturbation strategy for the machine unlearning process, the unlearning model and the retraining model essentially exhibit the characteristic of model indistinguishability. For datasets with i.i.d attributes, quantifying the unlearning effect by examining the accuracy on the remaining data $D_{RE}$ and the unlearning data $D_{UL}$ of the unlearning model can be challenging. As Fig.1 and Fig.6 show (accompanied with detailed discussions of results later), the accuracy of both $D_{RE}$ and $D_{UL}$ may be maintained at a fairly satisfactory level, the accuracy difference between $D_{RE}$ and $D_{UL}$ is quite small.
Figure 3: SPD-GAN architecture
To address the difficulty in measuring the effect of unlearning and further quantify the degree of forgetting caused by each unlearning strategies, we consider breaking the i.i.d attribute of $D_{UL}$.
Therefore, we design a SPD-GAN applying slight distributional perturbations to $D_{UL}$. It’s necessary to ensure that the added perturbation is minimal so that the model performance difference is not too large while also ensuring that the i.i.d attribute is indeed broken. Therefore, the objective is that perturbed unlearning data $D_{p}$ is guaranteed to perform well on the source model $M$, while the performance of the perturbed unlearning data $D_{p}$ on the unlearning model $M_{UL}$ differs significantly from that of the remaining data $D_{RE}$ on $M_{UL}$. That is to say:
$\centering P(M,D_{p})\approx P(M,D_{UL})~{}\wedge~{}P(M_{UL},D_{p})\ll P(M_{UL},D_{RE})\@add@centering$ (9)
Fig.3 shows the architecture of SPD-GAN, an autoencoder serves as the generator $G(\cdot)$ for noise generation. For the discriminator, the source model $M(\cdot)$ and the unlearning model $M_{UL}(\cdot)$ are employed as a joint discriminator.
To formally represent the aforementioned process, we define the following objective function. Initially, for the discriminator, $M(\cdot)$ and $M_{UL}(\cdot)$ are already trained, they are solely used to guide the training of the generator. In order to maximize the $M(\cdot)$’s accuracy on perturbed data $D_{p}$ while minimizing the $M_{UL}$’s accuracy on $D_{p}$, for generator, we have:
$\begin{split}L_{G}&=\mathbb{E}_{z\sim p_{z}(z)}[\log(1-M(G(z)+x))]\\ &\quad-\mathbb{E}_{z\sim p_{z}(z)}[\log M_{UL}(G(z)+x)]\end{split}$ (10)
Here, $\mathbb{E}$ denotes expectation, $z\sim p_{z}(z)$ denotes sampling from the latent variable distribution. Loss function are defined as:
$\centering loss=loss_{M(\cdot)}-\eta*loss_{M_{UL}(\cdot)}\@add@centering$ (11)
where $\eta$ is a coefficient to balance the $M$’s classification accuracy and $M_{UL}$’s unlearning effect. As Equation.1 explained, our goal is to explicitly demonstrate the degree of unlearning in terms of performance. Once completing the training of SPD-GAN, the performance difference between $M(\cdot)$ and $M_{UL}(\cdot)$ on perturbed data $D_{p}$ is indicative of unlearning degree. That is:
$\centering degree=P(M,D_{p})-P(M_{UL},D_{p})\@add@centering$ (12)
Here, we use ‘accuracy’ to characterize performance $P(\cdot,\cdot)$. In other words, we use $D_{p}$ as an approximation of $D_{UL}$, to measure the performance difference of $D_{UL}$ that does not conform to the i.i.d attribute on models before and after unlearning, thereby characterizing it as the degree of unlearning. Assuming total classes of data is $C$, the unlearning degree ranges in $[0,1-1/C]$.
Unlearning degree calculating process is detailed in Algorithm 2.
Algorithm 2 Unlearning Degree Evaluation
0:  Joint discriminators $M(\cdot)$ and $M_{UL}(\cdot)$, Generator $G(\cdot)$, coefficient $\eta$
1:  while Epochs do
2:     $\tilde{D_{UL}}=G(D_{UL})+D_{UL}$
3:     $y_{S}\leftarrow M(\tilde{D_{UL}})$
4:     $y_{U}\leftarrow M_{UL}(\tilde{D_{UL}})$
5:     $loss_{M(\cdot)}\leftarrow cross\_entropy(y_{S},Y)$
6:     $loss_{M_{UL}(\cdot)}\leftarrow cross\_entropy(y_{U},Y)$
7:     Minimize $loss\leftarrow loss_{M(\cdot)}-\eta*loss_{M_{UL}(\cdot)}$
8:  end while
9:  $D_{p}=G(D_{UL})+D_{UL}$
10:  $degree=P(M,D_{p})-P(M_{UL},D_{p})$
Figure 4: Visualization of images before and after SPD-GAN applied on ResNet18.
3.4 Time Complexity Analysis
Theorem 1
For each epoch, the computational time of $M_{UL}$ is less than that of the retraining model $M_{RE}$.
Proof: Let the time for a single forward pass and back-propagation of the retraining model $M_{RE}$ on the data set $D_{RE}$ be $\mathcal{O}(N)$, where $N$ denotes the model’s parameter count.
Considering $M_{UL}$, its loss function can be represented as: $L(\theta)=L(w)+\delta L(w)$ with $\delta L(w)$ representing the change due to data removal. For a Top-K unlearning strategy, the gradient update is expressed as $\delta w_{k}=-\alpha\cdot\frac{\partial\delta L}{\partial w_{k}}$ and $\delta w_{i}=0$ for $w_{i}\notin w_{k}$, yielding a time complexity of $\mathcal{O}(K)$ for gradient recomputation. Similarly, for the Random-k strategy, the complexity is $\mathcal{O}(k\cdot N)$ with $k\in(0,1)$. Hence, strategies based on partial parameters perturbed have significantly reduced epoch time compared to full retraining since both $\mathcal{O}(K)$ and $\mathcal{O}(k\cdot N)$ is smaller than $\mathcal{O}(N)$.
Theorem 2
$M_{UL}$ requires fewer epochs to converge compared to the retraining model $M_{RE}$.
Proof: Firstly, suppose $\theta$ is the parameter of the model after perturbation, and $\theta_{rand}$ is the parameter of random initialization, thus:
$\centering L(\theta;D_{RE})\leq L(\theta_{rand};D_{RE})\@add@centering$ (13)
Where $L(\theta,D_{RE})$ denotes the loss of the $D_{RE}$ under the parameter $\theta$. As indicated by Equation 13, the initial loss starting from $\theta$ is expected to be lower.
Furthermore, when considering the magnitude and direction of the gradient $\bigtriangledown L(\theta;D_{RE})$, since $\theta$ represents an optimized solution, for the majority of normalized directions $d$ (where its magnitude is standardized), we have:
$\centering d^{T}\bigtriangledown L(\theta;D_{RE})\leq d^{T}\bigtriangledown L(\theta_{rand};D_{RE})\@add@centering$ (14)
Starting from $\theta$, the requisite step size might be reduced, implying that the gradient’s direction remains more consistent.
Given that the gradient direction originating from $\theta$ exhibits stability and that the starting loss value is already minimized, the iterative optimization is poised to converge more swiftly. Suppose the complete retraining process requires $e_{1}$ epochs, while Top-K machine unlearning takes $e_{2}$ epochs, Random-k takes $e_{3}$ epochs. Consequently, we deduce that: $e_{1}\times\mathcal{O}(N)>e_{2}\times\mathcal{O}(K)$, here both $e_{1}$ surpasses $e_{2}$ and $\mathcal{O}(N)$ exceeds $\mathcal{O}(K)$. Same for Random-k, $e_{1}\times\mathcal{O}(N)>e_{3}\times\mathcal{O}(k\cdot N)$.
Time cost for Top-K parameters calculation
We use Equation 4 to approximate the sensitivity of each parameter. Since $|w|=N$, the time complexity for both the forward pass and back-propagation is linear in the number of parameters, the overall time complexity for evaluating sensitivity remains $\mathcal{O}(N)$. Once the model structure is determined, the parameters with Top-K sensitivity of a specific model should be calculated only once.
4 Experiments
4.1 Baseline Methods
The impact of the first few layers on the model is mainly to extract the features of the data, and the latter few layers use these features to perform tasks. Following the setting of DBLP:journals/corr/abs-2201-06640 , we set $K=\{5,10\}$, use EU-K and CF-K as two of our baseline methods.
EU-K: Performing exact unlearning involves targeting the final $K$ layers of the model. In this process, the model parameters are held constant except the last $K$ layers. These last $K$ layers are then reinitialized and trained from the beginning.
CF-K: Catastrophic forgetting the last $K$ layers of the model. Differing from EU-K, there’s no need to reinitialize the parameters of the last $K$ layers. The parameters of all layers except the last $K$ are held constant.
Retrain: Retrain the model using $D_{RE}$ from scratch.
4.2 Evaluation Metrics
To adequately assess the performance of machine unlearning, our considerations are twofold: the effectiveness of the unlearning process and the model generalization. Therefore, we propose novel metrics, namely the forgetting rate and memory retention rate. At the same time, we use acceleration ratio to show the speedup and similarity to measure model indistinguishability.
Forgetting Rate (FR)
The forgetting rate is a measure of performance decay on the unlearning data, post the model’s unlearning phase. In essence, the forgetting rate quantifies the extent of data ‘forgotten’ by the model:
$FR=\frac{Acc_{before}-Acc_{after}}{Acc_{before}}$ (15)
Here, $Acc_{before}$ represents the model’s accuracy on the unlearning data prior to the unlearning process, and $Acc_{after}$ indicates the model’s accuracy on the same set, post-unlearning.
Memory Retention Rate (MRR)
In machine unlearning, it’s also crucial to assess the performance of the unlearning model on the remaining data. To this end, we employ the memory retention rate:
$MRR=\frac{Acc^{\prime}_{after}}{Acc_{before}}$ (16)
Here $Acc^{\prime}_{after}$ represents the accuracy of the unlearning model on the remaining data.
Acceleration ratio
In the preceding text, we analyzed the time complexity of the machine unlearning strategy based on partial parameter perturbation. We utilize ‘unlearn time’ to measure the time taken for the unlearning model to reach convergence and calculate the acceleration ratio relative to the retraining method.
Similarity
To assess the degree of model indistinguishability, we employ the $similarity=1-JS~{}divergence$ as a metric to quantify the similarity between the unlearning model and the retraining model.
4.3 Experimental Results
4.3.1 Training settings
We conduct experiments on CIFAR-10 krizhevsky2009learning datasets using ResNet18 he2016deep , VGG simonyan2014very , GoogLeNet szegedy2015going , and DenseNet huang2017densely models. The training dataset is randomly divided into $D_{UL}$ and $D_{RE}$ with $D_{UL}$ comprising 5%, 10%, 15% and 20%. For Top-K, we set $K=45$ for ResNet18, and for the Random-k, $k=5\%$. The coefficient $\lambda$ in Equation.8 is set to 0.1. We use Adam as optimizer, for EU-K and CF-K, we set learning rate 1e-4. The generator architecture in SPD-GAN consists of (two Conv2d)-BatchNorm-Conv2d-BatchNorm layers for the encoder, followed by ConvTranspose2d-BatchNorm-ConvTranspose2d-Conv2d layers for the decoder. The LeakyReLU and ReLU activation functions are used for the encoder and decoder, respectively. Activation function Tanh is used to limit the output within [-1, 1]. The coefficient $\eta$ in Equation.11 is 0.03. All the experiments are conducted on a Nvidia A100 GPU.
4.3.2 Determination of K value for Top-K
(a) Unlearning data ratio of 5%
(b) Unlearning data ratio of 10%
(c) Unlearning data ratio of 15%
(d) Unlearning data ratio of 20%
Figure 5: Accuracy difference between $D_{RE}$ and $D_{UL}$ for Top-K under unlearning data ratio at (a) 5%, (b) 10%, (c) 15% and (d) 20%. The maximum accuracy difference and its corresponding K value is annotated. Such K value is identified as the optimal.
To investigate the optimal value of $K$ for Top-K unlearning strategy, we conduct experiments with different values of $K$ within the range of 0 to 50 when the unlearning data ratio varies from 5% to 20%. Following the above settings, we apply Top-K to four models over 100 epochs. Since machine unlearning aims to eliminate the influence of unlearning data while trying to not hurt the performance of remaining data, we calculate the accuracy difference between $Acc_{RE}$ and $Acc_{UL}$. As illustrated in Fig.5, the $K$ value corresponding to the maximum accuracy difference is identified as the optimal. Considering the optimal K values under four unlearning data ratios, we set $K=45$ for ResNet18 and DenseNet, $K=30$ for VGG, and $K=50$ for GoogLeNet.
As shown in Fig.5, although we have chosen the $K$ corresponding to the maximum accuracy difference, in actual experiments, the selection of $K$ values fluctuating within the range of 0 to 50 does not bring significant variability to the Top-K unlearning. For example, in Fig.5(a), for ResNet18, setting K=25 versus 45 does not make a substantial difference. Therefore, one can avoid additional effort and computation required to determine any optimal $K$.
4.3.3 Effectiveness of Top-K and Random-k
Fig.6 illustrates the unlearning accuracy $Acc_{UL}$ and remaining accuracy $Acc_{RE}$ of four different models trained on the CIFAR-10 using Top-K and Random-k over 50 epochs with unlearning ratio 20%. We employ $Acc_{RE}$ instead of $Acc_{test}$ for two reasons. Firstly, we aim to account for the accuracy on $D_{RE}$. Secondly, the testing data also incorporates information closely related to $D_{UL}$, leading to a decrease in $Acc_{test}$. Fig.1 further shows the accuracy after applying the Random-k on the CIFAR-10 for 50 epochs. The accuracy of $D_{RE}$ is significantly higher than that of $D_{UL}$. It can be observed that the application of the Top-K achieves a 95% $Acc_{RE}$ across the four models, indicating a strong memory retention capability. Meanwhile, the accuracy difference between unlearning model on the $D_{RE}$ and the $D_{UL}$ shows that both the Top-K and Random-k achieve the unlearning effect more or less.
(a) ResNet18
(b) VGG
(c) DenseNet
(d) GoogLeNet
Figure 6: Top-K/Random-k unlearning accuracy (%) on (a) ResNet18, (b) VGG, (c) DenseNet and (d) GoogLeNet.
Strategy ResNet18 VGG DenseNet GoogLeNet
Top-K 45 30 45 50
Random-k 0.558M 1.002M 0.348M 0.308M
EU-5/CF-5 2.37M $6$ 49162 10634
EU-10/CF-10 4.73M 4.73M 0.178M 0.265M
Table 1: The number of parameters perturbed in different models under various unlearning strategies.
Ratio Strategy $Acc_{UL}$ $Acc_{RE}$ FR MRR Similarity Time
5% Top-K 80.01 97.61 19.97 97.63 88.37 76.97
Random-k 76.56 90.07 23.44 90.07 78.77 135.43
EU-5 85.68 98.44 14.31 98.44 84.88 534.76
EU-10 85.88 99.02 14.12 99.02 69.42 662.34
CF-5 86.52 99.16 13.48 99.16 78.65 586.6
CF-10 86.64 99.54 13.36 99.54 99.98 607.24
Retrain 84.84 100 15.16 100 100 1190.15
10% Top-K 80.14 97.19 19.86 97.19 85.36 192.86
Random-k 81.48 96.52 18.52 96.52 91.03 249.11
EU-5 86.7 98.88 13.29 98.88 82.25 541.38
EU-10 84.88 98.12 15.12 98.12 90.4 417.87
CF-5 85.52 97.89 14.48 97.89 70.27 515.13
CF-10 86.26 98.91 13.73 98.91 91.37 664.39
Retrain 84.8 100 15.2 100 100 1040.08
15% Top-K 75.48 96.8 24.52 96.8 89.88 203.23
Random-k 79.85 98.09 20.15 98.09 83.59 335.87
EU-5 84.61 97.46 15.38 97.46 85.11 494.60
EU-10 86.52 98.75 13.48 98.75 91.38 627.94
CF-5 87.64 99.46 12.35 99.46 85.27 558.31
CF-10 86.05 98.79 13.94 98.79 91.75 685.18
Retrain 84.85 100 15.15 100 100 950.51
20% Top-K 75.46 97.07 24.54 97.07 86.37 425.01
Random-k 77.83 98.99 22.17 98.99 94.51 638.63
EU-5 84.48 97.91 15.51 97.9 71.77 636.77
EU-10 83.19 96.81 16.81 96.81 83.92 741.51
CF-5 85.86 98.14 14.14 98.14 81.76 778.28
CF-10 85.69 99.17 14.31 99.17 91.9 979.65
Retrain 85.24 100 14.76 100 100 1400.9
Table 2: The experimental results of the ResNet18 under different unlearning ratios of the CIFAR-10. Lower $Acc_{UL}(\%)$ and Time (s) and higher other metrics (%) indicate better performance.
Tab.2 provides a more comprehensive overview of the effects of different unlearning strategies at various unlearning ratios. Across different unlearning ratios, Top-K consistently exhibits the lowest $Acc_{UL}$ values (80.14%, 75.48%, and 75.46%), along with corresponding forgetting rates (19.86%, 24.52%, and 24.54%). This suggests that, relative to other strategies, Top-K achieves the most effective reduction of the impact of $D_{UL}$ on the model when the unlearning ratio belongs to 10%, 15%, 20%. However, for the 5% unlearning ratio under Random-k, its corresponding $Acc_{RE}$ and MRR of 90.07% noticeably lag behind the results of other strategies. We believe this disparity is due to the instability of the strategy that employs random parameter perturbation, leading to inconsistent unlearning effects. In the meanwhile, across both $Acc_{RE}$ and MRR, the performance of $M_{UL}$ from nearly all unlearning strategies is comparable to that of retraining models, achieving close to 100% accuracy. As for similarity, there isn’t a significant distinction among different strategies, but this does indicate that model indistinguishability is maintained across diverse strategies.
(a) ResNet18
(b) VGG
(c) DenseNet
(d) GoogLeNet
Figure 7: Acceleration compared to retraining under different unlearning ratio on (a) ResNet18, (b) VGG, (c) DenseNet and (d) GoogLeNet
As discussed earlier, we believe that employing machine unlearning strategies offers a faster way to eliminate the influence of unlearning data on models compared to the retraining approach. The metric ’Time’ in Tab.2 indicates the time required for various strategies to complete unlearning. All strategies achieve an accelerations, but Top-K and Random-k take less time. As illustrated in Tab.1, this acceleration is due to the fact that, for experiments on ResNet18, Top-K and Random-k perturb less parameters than others. To further illustrate the acceleration effect of unlearning strategies on different models compared to the retraining method, we also conduct machine unlearning tasks on VGG, DenseNet, and GoogLeNet. Fig.7 shows the acceleration ratio of each unlearning strategy compared to retraining under different models and unlearning ratios. The acceleration effect of Top-K and Random-k is better than other strategies. Taking Fig.7(a) as an example, when the unlearning ratio is 5%, Top-K can achieve an acceleration of more than 15x, while Random-k achieves an acceleration close to 9x. The acceleration from EU-K and CF-K methods is not significant. As the unlearning ratio increases, the acceleration of Top-K and Random-k slows down. This is because, as more data is to be forgotten, the degree of modification to the model deepens. In conclusion, we can ascertain that the Top-K achieves optimal unlearning performance with minimal parameter perturbed, maintains considerable memory retention capability, and achieves the fastest acceleration effect.
Figure 8: $Acc_{RE}$ (%) comparison using Random-k, Top-K and mixed Random-TopK under different unlearning data ratios in [5%, 10%, 15%, 20%] on (a-d) ResNet18, (e-h) VGG, (i-l) DenseNet and (m-p) GoogLeNet.
4.3.4 Mixed Top-K and Random-k
In Random-k, parameters are chosen randomly for perturbation, while in Top-K, parameters to be perturbed are carefully calculated. To investigate whether the random selection in Random-k is efficient enough or can be further improved, we designed a set of experiments in which $K$ in $k\%$ parameter selected by Random-k is replaced by those chosen by Top-K. This formed a mixed Random-TopK unlearning strategy, and related experiments are conducted using settings similar to the aforementioned ones. The $Acc_{RE}$ for Top-K, Random-k, and Random-TopK after 100 training epochs are shown in Fig.8.
Based on the $Acc_{RE}$ trends for Top-K, Random-K, and Random-TopK, at the beginning of training, the $Acc_{RE}$ for Top-K is higher than that for Random-K and Random-TopK. This implies that as the number of perturbed parameters in the model increases, more knowledge is eliminated. For Random-TopK, its $Acc_{RE}$ is lower than both Top-K and Random-K, indicating that over-modifying parameter based on those with the highest sensitivity will be more time-consuming. For models like ResNet18 (Fig.8(a)-8(d)), VGG (Fig.8(e)-8(h)), and DenseNet (Fig.8(i)-8(l)), Random-TopK still exhibits a certain upward trend in $Acc_{RE}$, suggesting that even with excessive parameter modifications, Random-TopK can still achieve partial unlearning effects. However, we notice that when using the Random-Topk stragegy on GoogLeNet (Fig.8(m)-8(p)), the model fell into a state of complete non-learning. We believe this may be due to the different sensitivities and parameter dependencies between the different branches of the Inception module in GoogLeNet. Employing the Random-TopK strategy might introduce complexities, adversely affecting certain branches.
4.3.5 Unlearning Degree Evaluation
Fig.4 shows the original images and the perturbed images when SPD-GAN is applied on ResNet18. When noise is added to $D_{UL}$, it doesn’t impact people’s ability to correctly recognize images. However, it does disrupt the $D_{UL}$ distribution.
Figure 9: Unlearning degree (%, ranges in [0, 90%]) evaluation for different unlearning strategies on ResNet18 using CIFAR-10 under different unlearning data ratio (5% - 20%). The higher unlearning degree is regarded as the better.
As aforementioned, in order to measure the unlearning degree for each unlearning strategy, we use Equation.12 to calculate the performance difference on the perturbed data $D_{p}$ generated by SPD-GAN between the source model $M$ and unlearning model $M_{UL}$. Utilizing source model $M$ and unlearning model $M_{UL}$ obtained from applying different unlearning strategies on CIFAR-10 with ResNet18 as the discriminator, the noise generator $G(\cdot)$ is trained.
The resulting unlearning degree is illustrated in Fig.9. The uniform random output accuracy for CIFAR-10 is around 10%, therefore, when we conduct experiments with CIFAR-10, the range of unlearning degree is [0, 90%]. From the experimental results in Fig.9, it can be observed that the use of Top-K achieves deepest unlearning degree that close to 90%. Random-k follows, achieving the second-deepest unlearning degree. However, due to EU-K and CF-K freezing most of parameters and modifying the last $K$ layers, their degree of unlearning is significantly lower than that of Top-K and Random-k. Moreover, at lower ratio of unlearning data, such as 5%, the degree of unlearning across different unlearning strategies if relatively higher.
5 Conclusion
In this paper, we propose fine-grained model weights perturbation methods to achieve inexact machine unlearning. Our method ensures a higher degree of unlearning the information from dataset $D_{UL}$ while maintaining model indistinguishability and achieving largest acceleration effect. Furthermore, after using SPD-GAN to break the i.i.d attribute of the unlearning data, we can quantify the degree of unlearning. The results further demonstrate that our proposed Top-K based inexact unlearning strategy achieves the best unlearning effect.
Limitations and Future Work
Although the Top-K strategy can achieve the best unlearning effect with the smallest perturbation ratio and the fastest acceleration, our current analysis of parameter sensitivity in the model ignores the dependence between parameters. One possible direction is to redesign and select Top-K parameters for unlearning process considering the dependencies between parameters. By examining these dependencies during the unlearning process, it would become possible to interpret and explain the effects of unlearning on different components of the model, enhancing interpretability and transparency. This, in turn, is expected to facilitate better optimization of the unlearning process to improve overall model performance.
Acknowledgments
The work is supported by the National Key Research and Development Program of China (2021ZD40303), the National Natural Science Foundation of China (Grant Nos. 62225205, 92055213), Natural Science Foundation of Hunan Province of China (2021JJ10023), Shenzhen Basic Research Project (Natural Science Foundation) (JCYJ20210324140002006).